OpenStudy (anonymous):

xavier is 3 years more than twice as old as yoko. the sum of their ages is 84. how old are they?

OpenStudy (anonymous):

x=2y+3 x+y=84 (2y+3)+y=84 3y+3=84 3y=81 y=27 x=2(27)+3 x+57

OpenStudy (anonymous):

* x = 57